OpenStudy (waheguru):

p(x)=-5x^2+400x-2550 repersents the profit made when x amount of money is spend on advertising How much money should be spend on advertising to have a profit of atleast 4 000 000

OpenStudy (waheguru):

Can someone explain how to solve this?

OpenStudy (waheguru):

I tried to set p to 4 000 000 but then what do I do?

OpenStudy (anonymous):

\[ -5x^2+400x-2550=4 000 0 \]

OpenStudy (anonymous):

Subtract \(4 000 0\) and use the quadratic equation.

OpenStudy (waheguru):

its 4 000 000

OpenStudy (waheguru):

It there another way other than the quadratic equation?

OpenStudy (waheguru):

Partal Factoring?

OpenStudy (anonymous):

You can complete the square.

OpenStudy (waheguru):

But then the x value stays 40? which is not right

OpenStudy (mertsj):

\[-5x^2+400x-2550=4000000\] \[5x^2-400x+2550=-4000000\] \[5x^2-400x+4002550=0\] \[x^2-80x+800510=0\]

OpenStudy (waheguru):

how did you get 800510?

OpenStudy (mertsj):

Did you notice that I divided both sides of the equation by 5?

OpenStudy (waheguru):

whats next?

OpenStudy (mertsj):

Something is wrong with your posting because that equation has only imaginary roots.

OpenStudy (waheguru):

its 4000

OpenStudy (waheguru):

Not 4000 000 the question said in thousands but then we make millions

OpenStudy (waheguru):

so its really 4000

OpenStudy (mertsj):

Well...that should make a significant difference.

OpenStudy (waheguru):

sorry

OpenStudy (mertsj):

That equation has real roots but it will not factor so plug it into the quadratic formula.

OpenStudy (waheguru):

so we have no choice but to use the quadracic formuula... the lesson we are learning is partial factoring i was guessing it would be related somehow?

OpenStudy (mertsj):

Yes. You could use partial factoring.

OpenStudy (waheguru):

Could you explain how?

OpenStudy (waheguru):

because I am always getting 40 as the x value

OpenStudy (mertsj):

-5((x-80)x+510)=4000

OpenStudy (waheguru):

-5x^2+400x-2550-4000 -5x^2+400x-6550 -5x(x-80)-6550 and then the x value is 40 80/2 - 40...

OpenStudy (waheguru):

What is wrong?

OpenStudy (mertsj):

I don't see how you get the x value of 40 from what you have posted.

OpenStudy (waheguru):

What should x be then?

OpenStudy (waheguru):

in partial factoring dont we divide p/2 to give us the x value for the vertex?

OpenStudy (waheguru):

Can you show how to find x? please

OpenStudy (waheguru):

how can this be solved by partial facotring...???

OpenStudy (mertsj):

Perhaps that is not what is meant by partial factoring. Let me google it and see.

OpenStudy (waheguru):

Ill ask my teacher then...its ok

OpenStudy (mertsj):

\[-5x^2+400x-2550=4000\] \[-5x2+400x-6550=0\] \[5x^2-400x+6550=0\] \[x^2-80x+1310=y\] Let y = 1310 x(x-80)=0 x=0, x=80

hartnn (hartnn):

partial factoring means bringing into vertex form a(x+h)^2+k

hartnn (hartnn):

for solving quadratics using graphs

OpenStudy (waheguru):

Ill confirm with my teacher on monday thanks for your help @Mertsj

OpenStudy (mertsj):

yw